Write an equation of the line that passes through (10,-4) and is perpendicular to the line whose equation is y = 2/5x - 8.
Answer by ewatrrr(24785)


Hi,
Note:the standard slope-intercept form for an equation of a line is: y = mx + b
where m is the slope and b the y-intercept.
Line perpendicular to line y = 2/5x - 8, has slope = -(5/2)
Equation of this line:
y = -(5/2)x + b using ordered pair Pt(10,-4) to solve for b
-4 = -25 = b
21 = b -25,25,-25,25,
y = -(5/2)x + 21